OpenStudy (anonymous):

The number of baseball games Bernie won this year was 25% less than the number of games he won last year. If Bernie won 24 baseball games last year, how many games did he win this year?

OpenStudy (danjs):

Last year he won 24 This year he won 25% less than last year

OpenStudy (danjs):

so this year he won 75% of 24 games 24*0.75 = 18

OpenStudy (anonymous):

Thanks DanJS👌

OpenStudy (mathstudent55):

Other way of calculating: This year hew won 25% fewer games than last year. 25% of 24 is 25% * 24 = 0.25 * 24 = 6 Last year he won 24, this year he won 24 - 6 = 18

OpenStudy (anonymous):

Thank you for another way to help me ✌ Mathstudent55

OpenStudy (mathstudent55):

You're welcome.